# Prompts to Proxies: Emulating Human Preferences via a Compact LLM Ensemble

### v1

@@ C.3 Example Endowments
They are a 32-year-old nonbinary software developer living in the suburbs. They highly value personal liberty and have a strong work ethic, often putting in extra hours for their projects. With a master’s degree in computer science, they frequently engage with scientific research, feeling that understanding complex topics fosters innovation. Their income is above average, and they embrace intelligence as a vital asset for personal growth.
An example of an endowment for a white male progressive in his 60s living in a small town
[…]
Pat is a 67-year-old White male from a small town. He identifies as a progressive and believes strongly in the importance of intelligence in community discussions and personal growth. Pat, who identifies as heterosexual, is skeptical of the increasing influence of big corporations on local politics, and he often seeks media that highlights grassroots movements and community activism. He balances his media consumption between traditional outlets and local storytelling platforms, knowing the value of good journalism.
Appendix D Extended Literature Review
D.1 Large Language Models to Emulate Human Preferences
@@
Leveraging the ability of LLMs to emulate human or human-group responses with considerable fidelity, LLM alignment or personalization Tseng et al. [2024] has been largely focused on emulating human responses using various techniques. These include learning from feedback or reward models Bai et al. [2022], Christiano et al. [2017], under distribution shifts Leike et al. [2018], Krueger et al. [2021], finetuning Tan et al. [2024], retrieval augmented generation Prahlad et al. [2025], representation learning Ren et al. [2024], and using human-generated Horton [2023], Castricato et al. [2025], Argyle et al. [2023] prompts. However, when diverse responses need to be generated from many users or user groups instead of individuals, the data and time resources required for alignment grow linearly with the number of LLMs that need to be aligned. To reduce alignment costs, some works instead rely on AI-generated Schuller et al. [2024], Salminen et al. [2024], Ge et al. [2025], Simmons [2023] prompts, or alignment to multiple user groups Zhao et al. [2024], Mondal et al. [2025], even entire populations Cao et al. [2025]. However, even the best models struggle with this task Cao et al. [2025]. Furthermore, responses from LLMs have been found to have less variation than responses from real surveys Bisbee et al. [2024], and LLMs have a tendency to respond in the middle category Wang et al. [2024]. Consequently, it is imperative to address these limitations and devise data- and time-efficient approaches that enable LLMs to more accurately and reliably serve as proxies for survey populations. Our work contributes to this goal by introducing an active learning–inspired alignment framework that dynamically constructs a diverse agent basis under resource constraints and reconstructs population-level preferences via regression-based aggregation, without relying on ground-truth demographic profiles.
D.2 Data Selection and Active Learning
@@
Traditional data selection methods aim to improve the data efficiency of supervised deep learning models by identifying the most informative or important training examples for generalization Coleman et al. [2020], Paul et al. [2021]. Selection via Proxy (SVP) Coleman et al. [2020] uses a lightweight proxy model to estimate sample utility at a fraction of the cost, while GraNd and EL2N scores Paul et al. [2021] identify important training examples early in training.
Similarly, deep active learning applies query strategies to iteratively select the most informative samples from a large pool of data to be added to the training dataset for retraining or fine-tuning Astorga et al. [2024], Hübotter et al. [2025], Li et al. [2025]. Active learning strategies balance exploitation, focusing on uncertain or high-impact samples, with exploration, seeking diverse and novel samples to broaden coverage.
@@
Appendix E Mathematical Formalization of the Preference Reconstruction Theory
E.1 Space Definitions and Mapping Framework
We define four key spaces for formalizaing our alignment framework:
  • Attribute space $\mathcal{B}\subset\mathbb{R}^{d}$: each $b\in\mathcal{B}$ is a structured configuration of interpretable attributes drawn from the attribute bank (e.g., ideology, altruism, personality). Attributes serve as control handles that steer LLM behavior.
  • Endowment space $\mathcal{E}$: each $e\in\mathcal{E}$ is a natural language prompt or agent description instantiated from an attribute configuration $b\in\mathcal{B}$. Endowments serve as the actual inputs to the LLM. We treat $\mathcal{E}$ as a structured set, not a Euclidean space.
  • Preference space $\mathcal{P}\subset\mathbb{R}^{m}$: each $p\in\mathcal{P}$ represents a latent vector encoding a behavioral inclination or preference profile. This space is unobservable but assumed to mediate how attributes shape responses.
  • Response space $\mathcal{R}\subset\mathbb{R}^{n}$: each $r\in\mathcal{R}$ is an observable vector of responses to survey questions (e.g., binary or Likert-coded).
Based on the four spaces, we define three mappings:
  • Attribute-to-Endowment Mapping ($\phi:\mathcal{B}\to\mathcal{E}$): This is a stochastic mapping that transforms a symbolic attribute configuration into a textual prompt or persona instantiation readily parsable by an LLM. Since multiple distinct prompts can be generated from the same attributes, we model $\phi$ as sampling from a conditional distribution over prompts given attributes.
  • Endowment-to-Preference Mapping ($g:\mathcal{E}\to\mathcal{P}$): This captures the internal LLM process by which a given prompt (endowment) induces a latent preference state. This mapping is governed by the LLM’s internal representations, shaped jointly by its large-scale pretraining corpora, supervised fine-tuning (SFT), reinforcement learning from human feedback (RLHF). Albeit unobservable, this preference vector mediates the model’s behavioral tendencies in downstream response generation.
  • Preference-to-Response Mapping ($h:\mathcal{P}\to\mathcal{R}$): This final mapping generates a survey response vector based on the latent preference state. It represents how an agent’s inclinations are expressed as answers to survey questions.
E.2 Preference Reconstruction via a Functional Basis
The overall agent behavior function is thus a composition of the three mappings:
$f=h\circ g\circ\phi:\mathcal{B}\rightarrow\mathcal{E}\rightarrow\mathcal{P}\rightarrow\mathcal{R}$
where $f(b)$ denotes the response behavior of an LLM agent endowed based on attributes $b$.
Given a set of sampled agents $\mathcal{A}=\{b_{1},\dots,b_{k}\}$, we obtain a response basis $\hat{\mathcal{R}}=\{f(b)\mid b\in\mathcal{A}\}$. We then treat alignment as a projection task:
$\min_{w\in\mathbb{R}^{|\mathcal{A}|}}\left\|\sum_{b\in\mathcal{A}}w_{b}f(b)-r_{\text{pop}}\right\|^{2}$ (8)
where $r_{\text{pop}}$ is the target aggregate response vector, and the weights $w_{b}$ reflect the inferred contribution of each proxy agent to the population preference.
E.3 Alignment as a Two-Stage Problem
We frame the alignment process as a two-stage problem grounded in mappings across attribute, endowment, preference and responses spaces.
Stage 1: Constructing a functional basis based on response variability.
The goal of the first stage is to construct a diverse and expressive set of LLM agents whose latent preferences collectively span the space where the ground-truth population preference resides. Since latent preferences are not directly observable, we use response variability as an indirect gauge, leveraging the preference-to-response mapping ($h:\mathcal{P}\to\mathcal{R}$). Starting from a sampled set of attribute configurations $\mathcal{A}=\{b_{1},\dots,b_{k}\}\subset\mathcal{B}$, we apply the composition mapping $f:\mathcal{B}\to\mathcal{R}$ to obtain a set of simulated agent responses:
$\hat{\mathcal{R}}=\{f(b)\mid b\in\mathcal{A}\}\subset\mathcal{R}.$
We interpret variability in $\hat{\mathcal{R}}$ as an indirect signal of the latent variation in $\hat{\mathcal{P}}$. The design objective is to ensure that the span of $\hat{\mathcal{R}}$ is sufficiently rich so that the latent preference vectors form a functional basis for reconstructing population-level behavior. To optimize this objective, we operate in the more interpretable and tractable attribute space $\mathcal{B}$ rather than the high-dimensional preference space $\mathcal{P}$. By actively exploring semantically meaningful regions of $\mathcal{B}$—analogous to domain-aware active learning strategies Hao et al. [2024] and latent-manifold transformations Jin et al. [2024]—we adaptively improve coverage of $\hat{\mathcal{P}}$.
Stage 2: Projecting onto population responses.
In the second stage, we assume access to a population-level response vector $r_{\text{pop}}\in\mathcal{R}$, such as aggregate survey results. The construction of the agent ensemble is done by solving the optimization problem defined by (8), where each weight $w_{b}$ reflects the inferred contribution of the agent defined by attribute set $b$ in reconstructing the aggregate signal.
Together, these two stages—constructing a coverage-maximizing basis and estimating population-aligned weights—enable efficient and demographically-agnostic preference reconstruction using steerable LLM agents.

### v2

@@ D.2.1 Experiment Setup
Table 7 reports results across the three locales. Each experiment is repeated 3 times.
D.2.2 Results and Discussion
Table 7: P2P performance on WVS Wave 7 (US, Great Britain, Hong Kong SAR) under a 7:1.5:1.5 question split, using gemini-2.0-flash as the backend. Reported are mean $\pm$ std over 3 repeated runs.
Metric WVS W7 US WVS W7 GB WVS W7 HK
Avg. Entropy $\uparrow$ 0.5720 $\pm$ 0.0149 0.5710 $\pm$ 0.0044 0.5478 $\pm$ 0.0211
Cost (USD) $\downarrow$ 1.7551 $\pm$ 0.0097 1.7095 $\pm$ 0.0031 1.7365 $\pm$ 0.0080
Test MSE (Lasso) $\downarrow$ 0.0149 $\pm$ 0.0005 0.0141 $\pm$ 0.0014 0.0242 $\pm$ 0.0013
Test 1-JSD (Lasso) $\uparrow$ 0.8030 $\pm$ 0.0055 0.7895 $\pm$ 0.0049 0.7590 $\pm$ 0.0046
Test MCPA (Lasso) $\uparrow$ 0.6667 $\pm$ 0.0152 0.6481 $\pm$ 0.0321 0.5789 $\pm$ 0.0263
Test EMD (Lasso) $\downarrow$ 0.1075 $\pm$ 0.0035 0.1038 $\pm$ 0.0081 0.1424 $\pm$ 0.0034
Test MSE (Elastic Net) $\downarrow$ 0.0147 $\pm$ 0.0005 0.0137 $\pm$ 0.0010 0.0244 $\pm$ 0.0005
Test 1-JSD (Elastic Net) $\uparrow$ 0.8066 $\pm$ 0.0067 0.7950 $\pm$ 0.0073 0.7602 $\pm$ 0.0032
Test MCPA (Elastic Net) $\uparrow$ 0.6930 $\pm$ 0.0152 0.6296 $\pm$ 0.0321 0.5877 $\pm$ 0.0304
Test EMD (Elastic Net) $\downarrow$ 0.1066 $\pm$ 0.0025 0.1004 $\pm$ 0.0054 0.1430 $\pm$ 0.0046
Table 8: Comparison of P2P performance on ATP (US) and WVS (US). ATP values are mean $\pm$ std over 14 waves; WVS values are mean $\pm$ std over 3 runs.
Metric ATP (14 waves) WVS US
Test MSE (Lasso) $\downarrow$ 0.014 $\pm$ 0.004 0.015 $\pm$ 0.001
Avg. Entropy $\uparrow$ 0.53 $\pm$ 0.06 0.57 $\pm$ 0.01
Cost (USD) $\downarrow$ 0.8 $\pm$ 0.2 1.76 $\pm$ 0.01
As shown in Table 7, performance is higher for US (MSE $\approx$ 0.015) and GB (MSE $\approx$ 0.014) and somewhat lower for HK (MSE $\approx$ 0.024), aligning with broader findings that large language models trained primarily on English and Western-centric data tend to better match Western public opinion than non-Western populations (Durmus et al., 2024). The HK performance gap motivates the error analysis presented in Section 4.4 of the main paper. Despite the gap, the HK results remain within the performance range observed across our 14-wave ATP study presented in the main paper (cf. Table 1), suggesting that while locale affects difficulty, P2P remains broadly applicable. Compared with ATP, WVS results register a higher per-survey cost at approximately \$1.73, due to the higher number of survey questions (185 for US and GB, 172 for HK). Notably, the US results are nearly identical across survey programs (ATP: 0.014; WVS: 0.015), providing evidence against program-specific overfitting (Table 8).
D.3 Ablation Studies
D.3.1 Attributes
As discussed in Section 3.2, attributes play the role of control handles in P2P’s generation logic, and freeform templates allow P2P to derive attributes from a specific survey or question, complementing the preset attribute bank with data-driven insights on endowment generation. To evaluation the potency of the freeform templates, we conduct an ablation in study in which we sequentially turn off question patching, survey patching and both. We additionally test the impact of per-endowment attribute cap through a setting where the default max attribute=10 is changed to 20.
Table 9: Attribute Ablation. Reported are mean $\pm$ std over 3 repeated runs.
Metric Avg. Entropy $\uparrow$ Test MSE $\downarrow$ Test MSE $\downarrow$ Cost (USD) $\downarrow$
(Lasso) (Enet)
Original 0.4451 $\pm$ .0256 0.0095 $\pm$ .0024 0.0079 $\pm$ .0026 0.9708 $\pm$ .0036
Survey patch off 0.4378 $\pm$ .0134 0.0103 $\pm$ .0011 0.0102 $\pm$ .0002 0.9785 $\pm$ .0122
Question patch off 0.3989 $\pm$ .0203 0.0114 $\pm$ .0016 0.0109 $\pm$ .0015 0.9698 $\pm$ .0011
Freeform off 0.3953 $\pm$ .0097 0.0128 $\pm$ .0008 0.0123 $\pm$ .0012 0.9688 $\pm$ .0015
max attributes = 20 0.4363 $\pm$ .0061 0.0124 $\pm$ .0046 0.0115 $\pm$ .0041 0.9880 $\pm$ .0167
Results in Table 9 suggest that performance of P2P is robust to attribute cap, whereas turning off freeform templates (especially question patching) degrades both average entropy and test MSE without significant cost gain. Therefore, from a performance vantage, it is recommended to keep the freeform templates for the extra benefits of data-driven insights.
D.3.2 Endowment Budget
We run P2P for 9 different endowment budgets, scaled linearly from 130 to 450. We increase the update steps of active endowment generation accordingly to ensure that at each step the number of new endowments generated is fixed at 20. We also include a budget of 110 with no updates as an ablation of adaptive sampling. Table 10 shows that on W42, a higher budget generally improve test performance but gains plateau after about three updates, indicating a moderate budget suffices.
Table 10: Performance of P2P with varying endowment budget. The average question entropy, test MSE (Lasso and ElasticNet), and generation cost are repeated. All values are mean $\pm$ std over 3 repeated runs.
(a) Budgets 110–210
Endowment Budget 110 130 170 210
Avg. Entropy $\uparrow$ 0.3777 $\pm$ 0.0191 0.3697 $\pm$ 0.0057 0.4180 $\pm$ 0.0080 0.4341 $\pm$ 0.0043
Test MSE (Lasso) $\downarrow$ 0.0144 $\pm$ 0.0018 0.0124 $\pm$ 0.0043 0.0167 $\pm$ 0.0041 0.0100 $\pm$ 0.0025
Test MSE (Enet) $\downarrow$ 0.0149 $\pm$ 0.0014 0.0123 $\pm$ 0.0045 0.0156 $\pm$ 0.0028 0.0103 $\pm$ 0.0026
Cost (USD) $\downarrow$ 0.3487 $\pm$ 0.0035 0.8043 $\pm$ 0.0032 0.4108 $\pm$ 0.0089 0.5453 $\pm$ 0.0063
(b) Budgets 250–330
Endowment Budget 250 290 330
Avg. Entropy $\uparrow$ 0.4179 $\pm$ 0.0165 0.4369 $\pm$ 0.0168 0.4308 $\pm$ 0.0208
Test MSE (Lasso) $\downarrow$ 0.0111 $\pm$ 0.0006 0.0090 $\pm$ 0.0013 0.0133 $\pm$ 0.0030
Test MSE (Enet) $\downarrow$ 0.0104 $\pm$ 0.0006 0.0097 $\pm$ 0.0017 0.0121 $\pm$ 0.0031
Cost (USD) $\downarrow$ 0.6715 $\pm$ 0.0129 0.9365 $\pm$ 0.0027 1.0688 $\pm$ 0.0119
(c) Budgets 370–450
Endowment Budget 370 410 450
Avg. Entropy $\uparrow$ 0.4216 $\pm$ 0.0179 0.4540 $\pm$ 0.0142 0.4433 $\pm$ 0.0092
Test MSE (Lasso) $\downarrow$ 0.0140 $\pm$ 0.0037 0.0080 $\pm$ 0.0006 0.0099 $\pm$ 0.0020
Test MSE (Enet) $\downarrow$ 0.0136 $\pm$ 0.0055 0.0084 $\pm$ 0.0004 0.0097 $\pm$ 0.0020
Cost (USD) $\downarrow$ 1.2050 $\pm$ 0.0179 1.3295 $\pm$ 0.0142 1.4619 $\pm$ 0.0092
D.3.3 Model Backend
We test P2P’s performance on five model backends: GPT-4.1-mini, GPT-4.1-nano, Gemini-2.0-flash, Gemini-2.5-flash and a locally hosted Qwen. Performance varies substantially, with the best average question entropy (0.471) and test MSE (0.009) achieved by GPT-4.1-mini (Table 11). Gemini-2.0-flash ranks second on these metrics but is markedly more cost-efficient, requiring only one quarter of the cost of GPT-4.1-mini. Smaller models, like GPT-4.1-nano and Gemini-2.5-flash-lite, perform worse, consistent with our theoretical intuition that limited model capacity constrains the learnability condition (Definition 2.2).
Table 11: Comparison of P2P performance on W42 using different model backends. Reported are average question entropy, test MSE (Lasso and ElasticNet), and generation cost. For each model, we fix the endowment budget at 300. All values are mean $\pm$ std over 3 repeated runs.
Model Avg. Entropy Test MSE (Lasso) Test MSE (Enet) Cost (USD)
GPT-4.1-mini 0.471 0.0090 0.0088 3.83
(0.017) (0.0014) (0.0014) (0.045)
GPT-4.1-nano 0.3996 0.0226 0.0213 0.9265
(0.0154) (0.0042) (0.0048) (0.0043)
Qwen 0.193 0.0204 0.0196
(0.040) (0.0014) (0.0027)
Gemini-2.0-flash 0.445 0.0104 0.0091 0.971
(0.026) (0.0038) (0.0046) (0.004)
Gemini-2.5-flash-lite 0.360 0.0149 0.0151 1.023
(0.0008) (0.0009) (0.0013) (0.0010)
Note. Qwen was run on a local server, thereby incurring no API cost.
D.3.4 Regression-based aggregation
To gauge the contribution of the regression-based aggregation module, we ablate it by replacing regression with simple averaging of agent responses, which increases the test MSE from 0.0104 to 0.0254 (Table 12). In essence, regression not only distills the agent collection into a more compact ensemble, but it also plays a crucial role in ensuring the population preference reflected in the survey data are actually learned by the ensemble.
Table 12: Comparison of P2P performance with and without the regression-based aggregation stage on ATP W42. In the ablation, agent responses from active endowment generation are combined by simple averaging. For the regression setting, we use test MSE for Lasso.
Metric Without regression With regression
Test MSE (Lasso) $\downarrow$ 0.0254 $\pm$ 0.00041 0.0104 $\pm$ 0.00383
We also compare the default option-level weighting against block weighting, where each question contributes equally to the loss. Performance remains comparable under both schemes for W42 (Table 13).
Table 13: Sensitivity check of weighting schemes. Reported are mean $\pm$ std over 3 repeated runs.
Weighting Train MSE $\downarrow$ Test MSE $\downarrow$ Train MSE $\downarrow$ Test MSE $\downarrow$
Scheme (Lasso) (Lasso) (Enet) (Enet)
Default 0.00526 $\pm$ .00141 0.00946 $\pm$ .00241 0.00353 $\pm$ .00161 0.00792 $\pm$ .00261
Block 0.00472 $\pm$ .00228 0.00884 $\pm$ .00281 0.00389 $\pm$ .00152 0.00797 $\pm$ .00263
D.4 Stress Test: Comparison with an SFT-Aligned Model under Topic Shift
P2P is designed as a plug-and-play preference alignment method that can be deployed on new populations with limited-data and limited-compute regimes. In the main paper, the baseline comparisons thus focus on prompting-based ensemble methods under the same constraints. As discussed in Appendix E, supervised finetuning (SFT) for a distribution-calibrated model offers an alternative approach under more lenient data and compute environments. An exemplar is Cao et al. (2025), where SFT is used to train a predictive model based on pooled World Values Survey (WVS) data from 46 countries, and conditioned on locale to directly predict answer distribution for a given survey question. They further test the SFT-aligned model on a distinct unseen set of WVS questions along with other cross-country and cross-dataset validation tasks.
In this regard, our work differs from Cao et al. (2025) not just in terms of training regimes, but also in key methodological premises. Specifically, we regard each survey panel as a benchmark for a specific population preference on a given topic at a given moment in time, where such preference is gaugeable through survey responses (cf. Figure 1, right panel). We do not assume that preferences on a topic domain can be reliably inferred if the training questions never probe that domain. For example, inferring political attitudes solely from non-political or religious items is a strong assumption for both humans and synthetic agents. Likewise, we do not assume preferences can carry over from one survey program to another, particularly when the surveys are conducted at different points in time and on different subpopulations of the same locale. Studying how aligned preferences generalize across topics, survey programs and over time is itself a key direction for future work, and will require both theoretical and empirical advances.
Nevertheless, Cao et al. (2025)’s setup constitutes an interesting stress test for P2P outside our studied scenarios. We present in this section a head-to-head comparison with their method on their WVS arena. We follow their protocol: Q1 and Q2 for train/validation and Q3 as test, which induces a topic shift from general attitudinal & religious/ethical items (Q1+Q2) to political interest and culture items (Q3). In our language, this asks: if we align to a population on broad attitudinal dimensions (Q1+Q2), how well does the aligned ensemble generalize to a new unseen dimension (Q3: political-cultural questions)? As in Section D.2, we focus on the three representative locales: US (overlaps ATP), Great Britain (Western Europe) and Hong Kong SAR (East Asia). For each, we filter Cao et al. (2025)’s data to that locale, train P2P (backend: gemini-2.0-flash; 300 endowments) on local Q1+Q2 (train+val), and evaluate on Q3. Table 14 shows the training schemes for P2P and Cao et al. (2025).
Table 14: Different training regimes for P2P and (Cao et al., 2025).
Setup Train+Val Entries Test Entries Model backend
(Q1 + Q2) (Q3)
P2P (US) 185 (150 + 35) 59 gemini-2.0-flash
P2P (GB) 172 (140 + 32) 60 gemini-2.0-flash
P2P (HK) 185 (150 + 35) 59 gemini-2.0-flash
Cao et al. (C1: 46 countries) 8427 (6841 + 1586) 2719 Llama3-8B-Instruct
Table 15: P2P vs. Cao et al. (2025) on WVS (Q3). Reported are mean $\pm$ std over 3 repeated runs.
Setup Method Test MSE $\downarrow$ Test 1-JSD $\uparrow$ Test MCPA $\uparrow$ Test EMD $\downarrow$
P2P (US) Lasso 0.0217 $\pm$ .0014 0.7492 $\pm$ .0036 0.6271 $\pm$ .0143 0.1188 $\pm$ .0048
P2P (US) Enet 0.0218 $\pm$ .0019 0.7497 $\pm$ .0051 0.6384 $\pm$ .0317 0.1184 $\pm$ .0066
P2P (GB) Lasso 0.0249 $\pm$ .0009 0.7416 $\pm$ .0055 0.4889 $\pm$ .0347 0.1296 $\pm$ .0045
P2P (GB) Enet 0.0249 $\pm$ .0009 0.7463 $\pm$ .0067 0.5111 $\pm$ .0096 0.1294 $\pm$ .0048
P2P (HK) Lasso 0.0284 $\pm$ .0023 0.7115 $\pm$ .0125 0.4802 $\pm$ .0353 0.1507 $\pm$ .0122
P2P (HK) Enet 0.0290 $\pm$ .0023 0.7151 $\pm$ .0133 0.4689 $\pm$ .0298 0.1532 $\pm$ .0108
Cao et al. SFT 0.777 0.43 *
* EMD not reported due to incompatible definition in Cao et al. (2025).
Table 15 presents the results on the performance comparison. In summary, P2P achieves 1-JSD on Q3 comparable to Cao et al. (2025)’s SFT model and matches or exceeds their majority-class prediction accuracy (MCPA), despite using far fewer training entries per locale and no fine-tuning (Table 14). Following our earlier discussion, we interpret performance on Q3 as evidence that some preference structure transfers across topics, rather than as a requirement that a preference-alignment method must perfectly extrapolate to arbitrary unseen domains.
Appendix E Extended Literature Review
E.1 Pluralistic Alignment
AI alignment refers to the process of ensuring that an AI system operates in accordance with human intentions and values, whether at an individual or aggregate level (Ji et al., 2023; Leike et al., 2018). As AI systems are adopted by increasingly diverse users, they must be designed to recognize and address a wider range of needs. This necessitates for pluralistic systems capable of capturing and representing the diversity of human values and perspectives (Sorensen et al., 2024b). Pluralistic alignment therefore seeks to align models not with a single gold standard, but with a diverse range of user preferences across various attributes.
Most works on pluralistic alignment touch on three complementary ways, namely overton, steerable, and distributional pluralism, in which a single AI model or system can support diversity of views. In overton pluralism, the model aims to output a whole spectrum of reasonable responses (Sorensen et al., 2024a). Methods to achieve this include looking at alignment shifts to transform distributions toward overton coverage (Lake et al., 2024) or few shot prompting from community-specialized LLMs (Feng et al., 2024). In steerable pluralism, models are typically steered to adopt or favor particular perspectives or value attributes, often through few-shot examples (Feng et al., 2024; Adams et al., 2025). By contrast, our approach steers purely through attributes, with the goal of constructing diverse agents that embody different perspectives. In distributional pluralism, the distribution of the model over possible answers is intended to match that of some target population (Sorensen et al., 2024b), often achieved through model fine-tuning on human responses (Cao et al., 2025).
Another way to categorize pluralistic alignment work is by the amount of data they rely on. While most approached in pluralistic alignment involve post-training, they range from resource-intensive approaches that fine-tune on large annotated corpora to lightweight methods that operate with minimal or no additional data. With sufficient data, specialized community models can be fine-tuned and subsequently combined using model merging techniques (Yuan et al., 2024), few-shot prompting from community model outputs (Feng et al., 2024), or federated averaging (Srewa et al., 2025) to scale with user diversity. If limited data are available, an alternative line of post-training approaches leverage in-context learning, where similarity-based retrieval (Adams et al., 2025) or group-informed retrieval (Chen et al., 2024) is performed to select few-shot examples that guide models toward pluralistic alignment. In the absence of detailed data from individual pluralistic users, a third post-training approach is inference-time conditioning. Inference-time conditioning uses structured prompts that encode sociodemographic traits or behavioral dispositions to condition an LLM to emulate the responses of a specific agent. These agents are also known as endowments, personas (Castricato et al., 2025) or Silicon Samples (Horton, 2023; Argyle et al., 2023) and are lightweight to deploy and data-efficient. As inference-time conditioning is cheap and fast, they enable mixture or vector-based formulations of pluralistic alignment, which combine agents via multi-objective rewards to avoid averaging out minority perspectives (Chen et al., 2025; Feng et al., 2025). Therefore, P2P follows this line of work for pluralistic alignment. In the first step, P2P performs inference-time conditioning to generate a sufficient set of agents that adequately span the preference space. Thereafter, in the second step, it combines agents to achieve a compact yet pluralistic set of preferences that represent the target survey population.
E.1.1 Attributes
Pluralistic alignment works that use inference-time conditioning often steer agents toward specific profiles using attributes that include moral and value dimensions (Adams et al., 2025), or broad morals, values, characteristics, and perspectives (Sorensen et al., 2024a). Attributes may also refer to demographic features, that are more tangible. Examples of these attributes include age, sex, education, income, and religion (Castricato et al., 2025). P2P uses the term attributes more broadly to encompass both moral and value dimensions and demographic features. Bundling these together allows P2P to define and condition LLM agents in a way that reflects both who they are and how they evaluate options.
Other related works on pluralistic alignment relate to the curation of datasets, which support research on pluralistic alignment (Sorensen et al., 2024a; Zhou et al., 2025).
E.2 Large Language Models to Emulate Human Preferences
@@
Leveraging the ability of LLMs to emulate human or human-group responses with considerable fidelity, LLM alignment or personalization (Tseng et al., 2024) has been largely focused on emulating human responses using various techniques. These include learning from feedback or reward models (Bai et al., 2022; Christiano et al., 2017), under distribution shifts (Leike et al., 2018; Krueger et al., 2021), finetuning (Tan et al., 2024), retrieval augmented generation (Prahlad et al., 2025), representation learning (Ren et al., 2024), and using human-generated (Horton, 2023; Castricato et al., 2025; Argyle et al., 2023) prompts. However, when diverse responses need to be generated from many users or user groups instead of individuals, the data and time resources required for alignment grow linearly with the number of LLMs that need to be aligned. To reduce alignment costs, some works instead rely on AI-generated (Schuller et al., 2024; Salminen et al., 2024; Ge et al., 2025; Simmons, 2023) prompts, or alignment to multiple user groups (Mondal et al., 2025; Zhao et al., 2024b), even entire populations (Cao et al., 2025). However, even the best models struggle with this task (Cao et al., 2025). Furthermore, responses from LLMs have been found to have less variation than responses from real surveys (Bisbee et al., 2024), and LLMs have a tendency to respond in the middle category (Wang et al., 2024). Consequently, it is imperative to address these limitations and devise data- and time-efficient approaches that enable LLMs to more accurately and reliably serve as proxies for survey populations. Our work contributes to this goal by introducing an active learning–inspired alignment framework that dynamically constructs a diverse agent basis under resource constraints and reconstructs population-level preferences via regression-based aggregation, without relying on ground-truth demographic profiles.
E.3 Data Selection and Active Learning
@@
Traditional data selection methods aim to improve the data efficiency of supervised deep learning models by identifying the most informative or important training examples for generalization (Coleman et al., 2020; Paul et al., 2021). Selection via Proxy (SVP) (Coleman et al., 2020) uses a lightweight proxy model to estimate sample utility at a fraction of the cost, while GraNd and EL2N scores (Paul et al., 2021) identify important training examples early in training.
Similarly, deep active learning applies query strategies to iteratively select the most informative samples from a large pool of data to be added to the training dataset for retraining or fine-tuning (Astorga et al., 2024; Hübotter et al., 2025; Li et al., 2025). Active learning strategies balance exploitation, focusing on uncertain or high-impact samples, with exploration, seeking diverse and novel samples to broaden coverage.
@@
Appendix F Limitations and Future Work
Question weighting and regression loss
By design, P2P converts all multiple choice questions into binary questions before fitting a regression model. For example, a question with five options will turn into five questions with binary options. While this design ensures that regression can run smoothly with heterogeneous question types, it steers the model to prioritize fitting questions with more answer options, e.g., a 10-point Likert scale. Put differently, information contained in questions with fewer answer options, e.g., Yes/No, gets diluted after the binarization. Whether or not this design is desirable depends on the user’s assumption of the relationship between preference signal and choice granularity. The current design implicitly posits a positive relationship between the two—i.e., the finer the scale, the better a signal on the latent preference. For other assumptions, question reweighting or format transformation need to be done before the regression step. P2P uses option-level weighting as the default, but also offers block weighting as an option.
Survey data in, text out
The current design of P2P limits its use case to labeling: Given question and options, the aligned agent ensemble expresses its preference by offering a probability distribution over the given options. A key future research direction is to extend the use case to text completion, e.g., drafting resolution, offering suggestion and answering freeform survey questions. How to conflate individual agent textual responses into a representative aggregate response remains an under-explored area of research. Auction design (cf. Dütting et al., 2024) can be a promising aggregation algorithm in this case.
P2P as a benchmark for model steerability
In this study our primary goal is understanding preference reconstruction theory and its implementation through the two-stage alignment framework. While our focus is on the design aspects of the system, we acknowledge that the backend model’s steerability is also a determinant of P2P’s empirical performance. The ablation study using different backends offers first insights on this aspect. It also indicates P2P’s potential of being transformed into a benchmark pipeline for the study of model steerability in pluralistic alignment.
Temporal Alignment and Non-stationary Preference
Any method trained at a single moment risks becoming stale as attitudes and beliefs shift. This also affects P2P when considering using a fitted ensemble in future surveys. Our view is that synthetic agents should be periodically re-anchored using more recent, easier-to-collect data (e.g., shorter or cheaper surveys) and then tested on more demanding items. Architecturally, P2P is set up for this. The endowment generation and regression stages can be rerun on new waves, and one can track how learned weights and entropy patterns evolve, or when reconstruction quality begins to deteriorate. While P2P provides the machinery, systematic studies of preference shift and alignment across time and topics remain open.
Short-cut Learning and Generalization
Overfitting to survey artifacts is a key concern for any survey-based evaluation. This concern is closely related to what the robustness literature calls shortcut learning or Clever Hans behavior, where models exploit superficial regularities of a dataset instead of learning the underlying concept (Geirhos et al., 2020; Lapuschkin et al., 2019). Our empirical evaluation across 14 ATP waves is designed to mitigate the risk that our findings hinge on idiosyncrasies of a single wave, and the additional WVS experiments further mitigate the risk of ATP-specific shortcutting. While these steps serve as initial evidence against program-specific shortcutting, it does not constitute a complete solution to the generalization problem. A full multi-survey study (e.g., across GlobalOpinionQA and other polling corpora) is beyond the scope of this paper, but P2P ’s modular pipeline readily supports this kind of analysis. Fundamentally, we see this as partly a survey design challenge: if we want synthetic agents to learn stable preferences rather than survey-specific shortcuts, future work may need survey instruments that are explicitly constructed with learnability and cross-topic generalization in mind. In the last paragraph of Appendix , we suggest a semantic question module for P2P to quantify the expressiveness of survey questions, which could serve as a parallel line of work to mitigate shortcutting on the model-design side. Taken together, these directions highlight an important and promising interdisciplinary research frontier in our pursuit of alignment science.